OpenStudy (anonymous):

what percent of 600 is 84? proportion

OpenStudy (anonymous):

so you've asked about 8 questions that are all basically identical, can i help you understand how to do this so you can try the next one on your own?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

14% 84/600 = x/100 600 times x = 84 times 100 600x = 8400 600x/600 = 8400/600 x = 14 Check your work by multiplying 600 times 0.14 = 84

OpenStudy (anonymous):

"percent" means "per 100", for example, 50 out of every 200 of something is like saying 25 out of every 100. does that make sense?

OpenStudy (anonymous):

Yea but these problems are different though, they all dont work like that

OpenStudy (anonymous):

they're all the same because percent means converting X out of every Y to some number out of every 100

OpenStudy (anonymous):

oh.. well i guess I dont know how to set up the proportions for each problem..

OpenStudy (anonymous):

so for example 1 out of every 2 is like 50 out of every 100. Once you have the number out of every 100 it's a percent. So for example let's say we wanted to determine what percent was 5 out of 30: \[5 \over 30} = {x \over 100} \]

OpenStudy (anonymous):

\[ {5 \over 30}= {x \over 100} \]

OpenStudy (anonymous):

do you understand that so far? 5 out of 30 is equal to some number, X, out of 100.

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

Yes? ok thank you I understand that problem but I still dont get how to set up my problems.